OpenStudy (anonymous):

add and simplify: (x/x-1)-(2x/x^2-1) a. x/x-1 b. x/x+1 c. x^2-x/x^2-1

OpenStudy (anonymous):

\[\huge {x \over x-1} - {2x \over x^2-1}\] is that the problem?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

In the denominator of the second term use: \[a^2 - b^2 = (a-b)(a+b)\]

OpenStudy (anonymous):

(x-1)(x+1)(x/x-1)-(x-1)(x+1)(2x/x^2-1)

OpenStudy (anonymous):

so do i multiply the whole denominator by (x-1)(x+1)?

OpenStudy (anonymous):

(x-1)(x+1) is LCD

OpenStudy (anonymous):

How you got (x-1)(x+1) tell me first..

OpenStudy (anonymous):

Are you guessing it or you have done it by using the formula that I gave above???

OpenStudy (anonymous):

i did it by the formula that you gave me

OpenStudy (anonymous):

Yes.. Do you know about LCD ??

OpenStudy (anonymous):

yes, I just wasn't sure what it would be for this one.

OpenStudy (anonymous):

Can you factor out what is common in both the terms??

OpenStudy (anonymous):

Are you getting what I said??? Can you take common factor from both the terms ?

OpenStudy (anonymous):

is it x/x+1 ? I just factored out what was common.

OpenStudy (anonymous):

I think x/(x-1) is common in both the terms right???

OpenStudy (anonymous):

yes, so I factored it out. is that correct? and then I was left with x/x+1

OpenStudy (anonymous):

Didn't you get (x-1)/(x+1) ??? Check it carefully..

OpenStudy (anonymous):

See I show you the steps: \[\frac{x}{x-1} - \frac{2x}{(x-1)(x+1)} \implies \frac{x}{x-1} (1 - \frac{2}{x+1}) \implies \frac{x}{x-1}(\frac{x+1 - 2}{x+1})\]

OpenStudy (anonymous):

\[\implies (\frac{x}{x-1}) \times (\frac{x-1}{x+1}) \implies ??\]

OpenStudy (anonymous):

Getting what I have done or not ? If not then which step is creating problem for you??

OpenStudy (anonymous):

I'm slightly confused as to why you multiplied x/x-1 and (1-2/x+1) in the very beginning steps..

OpenStudy (anonymous):

I have factored out x/(x-1) out of the terms by using distributive property.. Getting??

OpenStudy (anonymous):

ok I get it

OpenStudy (anonymous):

Sure no???

OpenStudy (anonymous):

I didn't see it get explained this way, so maybe this will offer some clarity. The formula for adding or subtracting two fractions without a common denominator is this: \[\huge {a \over b} \pm {c \over d} = {a*d \pm c*b \over b*d}\] where, a, b, c and d are just whatever is in your fractions.